OpenStudy (anonymous):

1+cos(18x)=_____. a) 2sin^2 (18x) b) 2cos^2 (9x) c) 2cos^2 (18x) d) 2sin^2 (9x)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

hi need help

OpenStudy (anonymous):

yes i need help

OpenStudy (anonymous):

i gave you a medal

OpenStudy (anonymous):

with what

jimthompson5910 (jim_thompson5910):

\[\Large 1 + \cos(18x) = 1 + \cos(2*9x)\] \[\Large 1 + \cos(18x) = 1 + 2\cos^2(9x) - 1\] \[\Large 1 + \cos(18x) = 2\cos^2(9x)\] On step 2, I used this identity \[\Large \cos(2x) = \cos^2(x) - 1\] see this link for more trig identities http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf specifically see page 2 and the "Double Angle Formulas" section

jimthompson5910 (jim_thompson5910):

oops I made a typo, I meant to say \[\Large \cos(2x) = 2\cos^2(x) - 1\]

OpenStudy (anonymous):

how do i use that equation to find the answer

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

someone please help

jimthompson5910 (jim_thompson5910):

I practically gave you the answer. Read through my steps again.

OpenStudy (anonymous):

oh duh.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

You use the identity to replace cos(2*9x) with 2*cos^2(9x) - 1 the "9x" is thought of as the theta in \[\Large \cos(2\theta) = 2\cos^2(\theta) - 1\]

jimthompson5910 (jim_thompson5910):

np